### v1

@@ 6.3. Well-posedness of tangent Navier–Stokes for small data
In Section 6.2 we proved the existence of solutions of the tangent Navier–Stokes system (6.1) for $g=0$ and any datum $\mathbf{f}\in(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}$ for $p\geq 2$ and dimension $d\in\{2,3,4\}$. In this section we complement such a study by showing that for any dimension $d\geq 2$ and sufficiently small data $(\mathbf{f},g)\in(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}\times\mathbf{L}^{p}_{\#}(\Gamma)$ (with $g$ not necessarily zero and for $p$ within a certain range to be specified later), there exists a unique solution of system (6.1). Before stating our well-posedness result for problem (6.1) we recall the following property satisfied by the variational formulation (4.8) for the tangent Stokes equations: Since said problem defines a continuous linear isomorphism from $\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$ onto $(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}\times\mathrm{L}^{p^{*}}_{\#}(\Gamma)$, A.2 (Banach–Nečas–Babuška theorem) guarantees the existence of a positive constant $C_{\operatorname{Stokes}}>0$ such that for each $(\mathbf{v},q)\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma)\times\mathrm{L}^{p^{*}}_{\#}(\Gamma)$
@@
and for each $(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$
@@
Let $\frac{4}{3}\leq p<\infty$ if $d=2$ or $\frac{d}{2}\leq p<\infty$ if $d\geq 3$, and let $p^{*}=\frac{p}{p-1}$ be its dual exponent. Assume $\Gamma$ is of class $C^{3}$.
@@
$\sup_{\textbf{0}\neq(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)}$ $\frac{\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{v}+\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u})\overline{\mathbf{u}}\cdot\mathbf{v}-\int_{\Gamma}\uppi\operatorname{div}_{\Gamma}\mathbf{v}-\int_{\Gamma}q\operatorname{div}_{\Gamma}\mathbf{u}}{\|\mathbf{u}\|_{1,p;\Gamma}+\|\uppi\|_{0,p;\Gamma}}$ (6.31)
@@
$\geq\frac{C_{\operatorname{Stokes}}}{2}\left(\|\mathbf{v}\|_{1,p^{*};\Gamma}+\|q\|_{0,p^{*};\Gamma}\right),\qquad\forall(\mathbf{v},q)\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma)\times\mathrm{L}^{p^{*}}_{\#}(\Gamma)$
and analogously
$\sup_{\textbf{0}\neq(\mathbf{v},q)\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma)\times\mathrm{L}^{p^{*}}_{\#}(\Gamma)}$ $\frac{\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{v}+\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u})\overline{\mathbf{u}}\cdot\mathbf{v}-\int_{\Gamma}\uppi\operatorname{div}_{\Gamma}\mathbf{v}-\int_{\Gamma}q\operatorname{div}_{\Gamma}\mathbf{u}}{\|\mathbf{v}\|_{1,p^{*};\Gamma}+\|q\|_{0,p^{*};\Gamma}}$ (6.32)
$\geq C_{\operatorname{Stokes}}\left(\|\mathbf{u}\|_{1,p;\Gamma}+\|\uppi\|_{0,p;\Gamma}\right)-C_{\operatorname{conv}}\|\overline{\mathbf{u}}\|_{1,p;\Gamma}\|\mathbf{u}\|_{1,p;\Gamma},$
$\geq(C_{\operatorname{Stokes}}-C_{\operatorname{conv}}\|\overline{\mathbf{u}}\|_{1,p;\Gamma})\left(\|\mathbf{u}\|_{1,p;\Gamma}+\|\uppi\|_{0,p;\Gamma}\right)$
$\geq(C_{\operatorname{Stokes}}-C_{\operatorname{conv}}R)\left(\|\mathbf{u}\|_{1,p;\Gamma}+\|\uppi\|_{0,p;\Gamma}\right)$
@@
Consequently, by A.2 (Banach–Nečas–Babuška theorem), we deduce that the linear problem (6.30) is well-posed. In particular, it satisfies
$\|\mathbf{u}\|_{1,p;\Gamma}+\|\uppi\|_{0,p;\Gamma}\leq\frac{2}{C_{\operatorname{Stokes}}}\sup_{\begin{subarray}{c}(\mathbf{v},q)\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma)\times\mathrm{L}^{p^{*}}_{\#}(\Gamma)\\ (\mathbf{v},q)\neq\textbf{0}\end{subarray}}\frac{\langle\mathbf{f},\mathbf{v}\rangle-\int_{\Gamma}g\,q}{\|\mathbf{v}\|_{1,p^{*};\Gamma}+\|q\|_{0,p^{*};\Gamma}}\leq\frac{2}{C_{\operatorname{Stokes}}}\left(\|\mathbf{f}\|_{(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}}+\|g\|_{0,p;\Gamma}\right).$ (6.33)
@@
6.4. Higher $\mathrm{L}^{p}$-based Sobolev regularity
In this section we investigate the pick-up $\mathrm{L}^{p}$-based regularity of solutions to the tangent Navier–Stokes equations beyond the basic regularity $(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$ allowed by data $(\mathbf{f},g)$ smoother than in $(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}\times\mathrm{L}^{p}_{\#}(\Gamma)$. The results are consistent with those in Section 4.4 for the tangent Stokes system.
@@ 6.4. Higher $\mathrm{L}^{p}$-based Sobolev regularity
Let $\frac{4}{3}<p<\infty$ if $d=2$ or $\frac{d}{2}<p<\infty$ if $d\geq 3$ and let $p^{*}=\frac{p}{p-1}$ be its dual exponent. Moreover, assume $\Gamma$ is of class $C^{m+3}$ for some nonnegative integer $m$. If $(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$ is a solution of (6.1) with $\mathbf{f}\in\mathbf{W}^{m,p}_{t}(\Gamma)$ and $g\in\mathrm{W}^{m+1,p}_{\#}(\Gamma)$, then $(\mathbf{u},\uppi)$ belongs to the space $\mathbf{W}^{m+2,p}_{t}(\Gamma)\times\mathrm{W}^{m+1,p}_{\#}(\Gamma)$. Moreover, the following estimate holds
$\|\mathbf{u}\|_{m+2,p;\Gamma}+\|\uppi\|_{m+1,p;\Gamma}\leq F_{m,p}(\|\mathbf{f}\|_{m,p,\Gamma},\|g\|_{m+1,p;\Gamma},\|\mathbf{u}\|_{1,p;\Gamma}),$ (6.34)
@@
$\|\mathbf{u}\|_{m+2,p;\Gamma}+\|\uppi\|_{m+1,p;\Gamma}\leq F_{m,p}(\|\mathbf{f}\|_{m,p,\Gamma},\|g\|_{m+1,p;\Gamma},\|\mathbf{u}\|_{m+1,p;\Gamma}),$ (6.35)
from which (6.34) easily follows from successive nesting.
We first prove (6.35) the case $m=0$. For general $m$ the proof will follow by induction.
Step 1: Case $m=0$. Notice that $(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$ is a solution of the (weak form of) the tangent Stokes equations (4.8) with data $\overline{\mathbf{f}}:=\mathbf{f}-(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\in(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}$ and $\overline{g}:=g\in\mathrm{W}^{1,p}_{\#}(\Gamma)$. Hence, we will prove that $(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\in\mathbf{L}^{p}_{t}(\Gamma)$ and
$\|(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\|_{0,p;\Gamma}\leq\widetilde{F}_{p}(\|\mathbf{f}\|_{0,p;\Gamma},\|g\|_{1,p;\Gamma},\|\mathbf{u}\|_{1,p;\Gamma}),$ (6.36)
from which the assertion for $m=0$ follows from a direct application of 4.9 (higher $\mathrm{L}^{p}$-based regularity of the tangent Stokes equations). We will simply show the case $d\geq 3$ (the case $d=2$ is easier), for which there are three subcases: $p>d$, $p=d$ and $\frac{d}{2}<p<d$.
@@
    If $p>d$, then by 2.7 (Morrey’s inequality), we deduce that $\mathbf{u}\in\mathbf{C}(\Gamma)$, and hence $(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\in\mathbf{L}^{p}_{t}(\Gamma)$ with $\|(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\|_{0,p;\Gamma}\leq\|\nabla_{\Gamma}\mathbf{u}\|_{0,p;\Gamma}\|\mathbf{u}\|_{0,\infty;\Gamma}\leq C\|\mathbf{u}\|_{1,p;\Gamma}^{2}$.
@@
    If $p=d$, then from 2.6 (Gagliardo–Nirenber–Sobolev inequality), we deduce that $\mathbf{u}\in\mathbf{L}^{q}_{t}(\Gamma)$ for each $1<q<\infty$, and hence by Hölder’s inequality, $(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\in\mathbf{L}^{q}_{t}(\Gamma)$ for each $1<q<p=d$. By 4.9, we deduce that $(\mathbf{u},\uppi)\in\mathbf{W}^{2,q}_{t}(\Gamma)\times\mathrm{W}^{1,q}_{\#}(\Gamma)$ for each $1<q<d$. Choosing $q\in(\frac{d}{2},d)$, we further obtain by 2.6 and 2.7 that $\nabla_{\Gamma}\mathbf{u}\in\mathbb{L}^{\frac{dq}{d-q}}(\Gamma)$ and $\mathbf{u}\in\mathbf{C}(\Gamma)$; whence $(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\in\mathbf{L}^{\frac{dq}{d-q}}_{t}(\Gamma)\subseteq\mathbf{L}^{d}_{t}(\Gamma)$ (because $\frac{dq}{d-q}\geq d=p$) and, altogether,
    $\|(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\|_{0,d;\Gamma}\leq\|\nabla_{\Gamma}\mathbf{u}\|_{0,d;\Gamma}\|\mathbf{u}\|_{0,\infty;\Gamma}\lesssim\|\mathbf{u}\|_{1,d;\Gamma}\|\mathbf{u}\|_{2,q;\Gamma}\lesssim\|\mathbf{u}\|_{1,d;\Gamma}\left(\|\mathbf{f}\|_{0,q;\Gamma}+\|g\|_{1,q;\Gamma}+\|(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\|_{0,q;\Gamma}\right)\\ \leq\|\mathbf{u}\|_{1,d;\Gamma}\left(\|\mathbf{f}\|_{0,d;\Gamma}+\|g\|_{1,d;\Gamma}+\|\mathbf{u}\|_{1,d;\Gamma}\|\mathbf{u}\|_{0,\frac{dq}{d-q};\Gamma}\right)\lesssim\|\mathbf{u}\|_{1,d;\Gamma}\left(\|\mathbf{f}\|_{0,d;\Gamma}+\|g\|_{1,d;\Gamma}+\|\mathbf{u}\|_{1,d;\Gamma}^{2}\right).$
@@
    If $\frac{d}{2}<p<d$, then from 2.6, we deduce $\mathbf{u}\in\mathbf{L}^{\frac{dp}{d-p}}_{t}(\Gamma)$, and hence by Hölder’s inequality, $(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\in\mathbf{L}^{q_{0}}_{t}(\Gamma)$ and $\|(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\|_{0,q_{0};\Gamma}\leq C\|\mathbf{u}\|_{1,p;\Gamma}^{2}$, where $q_{0}:=\frac{dp}{2d-p}\in(\frac{d}{3},p)$. Then, by 4.9, we deduce that $(\mathbf{u},\uppi)\in\mathbf{W}^{2,q_{0}}_{t}(\Gamma)\times\mathrm{W}^{1,q_{0}}_{\#}(\Gamma)$.
    If $q_{0}<\frac{d}{2}$, then from 2.6 we deduce that $\nabla_{\Gamma}\mathbf{u}\in\mathbb{L}^{\frac{dq_{0}}{d-q_{0}}}(\Gamma)$ and $\mathbf{u}\in\mathbf{L}^{\frac{dq_{0}}{d-2q_{0}}}_{t}(\Gamma)$, and hence by Hölder’s inequality $(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\in\mathbf{L}^{q_{1}}_{t}(\Gamma)$ for $q_{1}:=\frac{dq_{0}}{2d-3q_{0}}$ and so again by 4.9, $(\mathbf{u},\uppi)\in\mathbf{W}^{2,\min\{p,q_{1}\}}_{t}(\Gamma)\times\mathrm{W}^{1,\min\{p,q_{1}\}}_{\#}(\Gamma)$. We can iterate this procedure as follows: if $q_{n}<\frac{d}{2}$, define
    $q_{n+1}:=\frac{dq_{n}}{2d-3q_{n}};$ (6.37)
    then $(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\in\mathbf{L}^{q_{n+1}}(\Gamma)$, $(\mathbf{u},\uppi)\in\mathbf{W}^{2,r_{n+1}}_{t}(\Gamma)\times\mathrm{W}^{1,r_{n+1}}_{\#}(\Gamma)$, with $r_{n+1}:=\min\{p,q_{n+1}\}$ and

### v2

@@ 5.3. Well-posedness of tangent Navier–Stokes for small data
In Section 5.2 we proved the existence of solutions of the tangent Navier–Stokes system (5.1) for $g=0$ and any datum $\mathbf{f}\in(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}$ for $p\in[2,\infty)$ and dimension $d\in\{2,3,4\}$. In this section we complement such a study by showing that for any dimension $d\geq 2$ and sufficiently small data $(\mathbf{f},g)\in(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}\times\mathbf{L}^{p}_{\#}(\Gamma)$ (with $g$ not necessarily zero and for $p$ within the range given by (5.2)), there exists a unique solution of system (5.1). Before stating our well-posedness result for problem (5.1) we recall the following property satisfied by the variational formulation (4.1) for the tangent Stokes equations: since said problem defines a continuous linear isomorphism from $\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$ onto $(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}\times\mathrm{L}^{p^{*}}_{\#}(\Gamma)$, [5, Theorem A.2] (Banach–Nečas–Babuška theorem) guarantees the existence of a positive constant $C_{\operatorname{Stokes}}>0$ such that for each $(\mathbf{v},q)\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma)\times\mathrm{L}^{p^{*}}_{\#}(\Gamma)$,
@@
and for each $(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$,
@@
Let $p\in[\frac{4}{3},\infty)$ if $d=2$ or $p\in[\frac{d}{2},\infty)$ if $d\geq 3$. Assume $\Gamma$ is of class $C^{2}$.
@@
$\sup_{\textbf{0}\neq(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)}$ $\frac{\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{v}+\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u})\overline{\mathbf{u}}\cdot\mathbf{v}-\int_{\Gamma}\uppi\operatorname{div}_{\Gamma}\mathbf{v}-\int_{\Gamma}q\operatorname{div}_{\Gamma}\mathbf{u}}{\|\mathbf{u}\|_{1,p;\Gamma}+\|\uppi\|_{0,p;\Gamma}}$
@@
$\geq\frac{C_{\operatorname{Stokes}}}{2}\left(\|\mathbf{v}\|_{1,p^{*};\Gamma}+\|q\|_{0,p^{*};\Gamma}\right),\qquad\forall(\mathbf{v},q)\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma)\times\mathrm{L}^{p^{*}}_{\#}(\Gamma),$
because $R<\frac{C_{\operatorname{Stokes}}}{2C_{\operatorname{conv}}}$. Analogously,
$\sup_{\textbf{0}\neq(\mathbf{v},q)\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma)\times\mathrm{L}^{p^{*}}_{\#}(\Gamma)}$ $\frac{\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{v}+\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u})\overline{\mathbf{u}}\cdot\mathbf{v}-\int_{\Gamma}\uppi\operatorname{div}_{\Gamma}\mathbf{v}-\int_{\Gamma}q\operatorname{div}_{\Gamma}\mathbf{u}}{\|\mathbf{v}\|_{1,p^{*};\Gamma}+\|q\|_{0,p^{*};\Gamma}}$
@@
Consequently, by [5, Theorem A.2] (Banach–Nečas–Babuška theorem), we deduce that the linear problem (5.30) is well-posed. In particular, it satisfies
$\|\mathbf{u}\|_{1,p;\Gamma}+\|\uppi\|_{0,p;\Gamma}\leq\frac{2}{C_{\operatorname{Stokes}}}\sup_{\begin{subarray}{c}(\mathbf{v},q)\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma)\times\mathrm{L}^{p^{*}}_{\#}(\Gamma)\\ (\mathbf{v},q)\neq\textbf{0}\end{subarray}}\frac{\langle\mathbf{f},\mathbf{v}\rangle-\int_{\Gamma}g\,q}{\|\mathbf{v}\|_{1,p^{*};\Gamma}+\|q\|_{0,p^{*};\Gamma}}\leq\frac{2}{C_{\operatorname{Stokes}}}\left(\|\mathbf{f}\|_{(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}}+\|g\|_{0,p;\Gamma}\right).$
@@
5.4. Higher $\mathrm{L}^{p}$-based Sobolev regularity of tangent Navier–Stokes
In this section we investigate the pick-up $\mathrm{L}^{p}$-based regularity of solutions to the tangent Navier–Stokes equations beyond the basic regularity $(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$ allowed by data $(\mathbf{f},g)$ smoother than in $(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}\times\mathrm{L}^{p}_{\#}(\Gamma)$.
The results are consistent with those in Section 4.3 for the tangent Stokes system.
@@ 5.4. Higher $\mathrm{L}^{p}$-based Sobolev regularity of tangent Navier–Stokes
Let $p\in(\frac{4}{3},\infty)$ if $d=2$ or $p\in(\frac{d}{2},\infty)$ if $d\geq 3$. Moreover, assume $\Gamma$ is of class $C^{m+2,1}$ for some nonnegative integer $m$. If $(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$ is a solution of (5.1) with $\mathbf{f}\in\mathbf{W}^{m,p}_{t}(\Gamma)$ and $g\in\mathrm{W}^{m+1,p}_{\#}(\Gamma)$, then $(\mathbf{u},\uppi)$ belongs to the space $\mathbf{W}^{m+2,p}_{t}(\Gamma)\times\mathrm{W}^{m+1,p}_{\#}(\Gamma)$. Moreover, the following estimate holds
$\|\mathbf{u}\|_{m+2,p;\Gamma}+\|\uppi\|_{m+1,p;\Gamma}\leq F_{m,p}\big(\|\mathbf{f}\|_{m,p;\Gamma},\|g\|_{m+1,p;\Gamma},\|\mathbf{u}\|_{1,p;\Gamma}\big),$ (5.31)
@@
$\|\mathbf{u}\|_{m+2,p;\Gamma}+\|\uppi\|_{m+1,p;\Gamma}\leq F_{m,p}\big(\|\mathbf{f}\|_{m,p;\Gamma},\|g\|_{m+1,p;\Gamma},\|\mathbf{u}\|_{m+1,p;\Gamma}\big),$ (5.32)
from which (5.31) easily follows from successive nesting. The proof hinges on successive applications of 4.8 (higher $\mathrm{L}^{p}$-based regularity of the tangent Stokes equations), 2.8 (Gagliardo–Nirenberg–Sobolev inequality) and 2.9 (Morrey’s inequality).
We first prove (5.32) for $m=0$. For general $m$ the proof will follow by induction.
Step 1: Case $m=0$. Notice that $(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$ is a solution of the (weak form of) the tangent Stokes equations (4.1) with data $\overline{\mathbf{f}}:=\mathbf{f}-(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\in(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}$ (cf. (5.3)) and $\overline{g}:=g\in\mathrm{W}^{1,p}_{\#}(\Gamma)$. Hence, we will prove that $(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\in\mathbf{L}^{p}_{t}(\Gamma)$ and
$\|(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\|_{0,p;\Gamma}\leq\widetilde{F}_{p}\big(\|\mathbf{f}\|_{0,p;\Gamma},\|g\|_{1,p;\Gamma},\|\mathbf{u}\|_{1,p;\Gamma}\big),$ (5.33)
from which the assertion for $m=0$ follows from a direct application of 4.8. We will simply show the case $d\geq 3$ (the case $d=2$ is easier), for which there are three subcases: $p>d$, $p=d$ and $\frac{d}{2}<p<d$.
@@
    If $p>d$, then by 2.9, we deduce that $\mathbf{u}\in\mathbf{C}(\Gamma)$, and hence $(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\in\mathbf{L}^{p}_{t}(\Gamma)$ with $\|(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\|_{0,p;\Gamma}\leq\|\nabla_{\Gamma}\mathbf{u}\|_{0,p;\Gamma}\|\mathbf{u}\|_{0,\infty;\Gamma}\leq C\|\mathbf{u}\|_{1,p;\Gamma}^{2}$.
@@
    If $p=d$, then from 2.8, we deduce that $\mathbf{u}\in\mathbf{L}^{q}_{t}(\Gamma)$ for each $q\in(1,\infty)$, and hence by Hölder’s inequality, $(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\in\mathbf{L}^{q}_{t}(\Gamma)$ for each $q\in(1,p)=(1,d)$. By 4.8, we infer that $(\mathbf{u},\uppi)\in\mathbf{W}^{2,q}_{t}(\Gamma)\times\mathrm{W}^{1,q}_{\#}(\Gamma)$ for each $q\in(1,d)$. Choosing $q\in(\frac{d}{2},d)$, we further obtain by 2.8 and 2.9 that $\nabla_{\Gamma}\mathbf{u}\in\mathbb{L}^{\frac{dq}{d-q}}(\Gamma)$ and $\mathbf{u}\in\mathbf{C}(\Gamma)$, with the estimates
    $\|\mathbf{u}\|_{0,\infty;\Gamma}\lesssim\|\mathbf{u}\|_{2,q;\Gamma}\lesssim\|\mathbf{f}\|_{0,q,\Gamma}+\|g\|_{1,q;\Gamma}+\|(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\|_{0,q,\Gamma}$
    and
    $\|(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\|_{0,q,\Gamma}\leq\|\nabla_{\Gamma}\mathbf{u}\|_{0,d;\Gamma}\|\mathbf{u}\|_{0,\frac{dq}{d-q};\Gamma}\lesssim\|\mathbf{u}\|_{1,d;\Gamma}^{2}.$
    This implies $(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\in\mathbf{L}^{\frac{dq}{d-q}}_{t}(\Gamma)\subseteq\mathbf{L}^{d}_{t}(\Gamma)$ (because $\frac{dq}{d-q}\geq d=p$) along with
    $\|(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\|_{0,d;\Gamma}\leq\|\nabla_{\Gamma}\mathbf{u}\|_{0,d;\Gamma}\|\mathbf{u}\|_{0,\infty;\Gamma}\lesssim\|\mathbf{u}\|_{1,d;\Gamma}\left(\|\mathbf{f}\|_{0,d;\Gamma}+\|g\|_{1,d;\Gamma}+\|\mathbf{u}\|_{1,d;\Gamma}^{2}\right).$
@@
    If $p\in(\frac{d}{2},d)$, then from 2.8, we deduce $\mathbf{u}\in\mathbf{L}^{\frac{dp}{d-p}}_{t}(\Gamma)$, and hence by Hölder’s inequality, $(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\in\mathbf{L}^{q_{0}}_{t}(\Gamma)$ and $\|(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\|_{0,q_{0};\Gamma}\lesssim\|\mathbf{u}\|_{1,p;\Gamma}^{2}$, where $q_{0}:=\frac{dp}{2d-p}\in(\frac{d}{3},p)$. Then, by 4.8, we deduce that $(\mathbf{u},\uppi)\in\mathbf{W}^{2,q_{0}}_{t}(\Gamma)\times\mathrm{W}^{1,q_{0}}_{\#}(\Gamma)$.
    If $q_{0}<\frac{d}{2}$, then from 2.8 we deduce that $\nabla_{\Gamma}\mathbf{u}\in\mathbb{L}^{\frac{dq_{0}}{d-q_{0}}}(\Gamma)$ and $\mathbf{u}\in\mathbf{L}^{\frac{dq_{0}}{d-2q_{0}}}_{t}(\Gamma)$, whence by Hölder’s inequality $(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\in\mathbf{L}^{q_{1}}_{t}(\Gamma)$ for $q_{1}:=\frac{dq_{0}}{2d-3q_{0}}$. Invoking again 4.8 yields $(\mathbf{u},\uppi)\in\mathbf{W}^{2,\min\{p,q_{1}\}}_{t}(\Gamma)\times\mathrm{W}^{1,\min\{p,q_{1}\}}_{\#}(\Gamma)$. We can iterate this procedure as follows: if $q_{n}<\frac{d}{2}<p$, define
    $q_{n+1}:=\frac{dq_{n}}{2d-3q_{n}}.$ (5.34)
    Then, if $(\mathbf{u},\uppi)\in\mathbf{W}^{2,q_{n}}_{t}(\Gamma)\times\mathrm{W}^{1,q_{n}}_{\#}(\Gamma)$, Hölder’s inequality and 2.8 imply that $(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\in\mathbf{L}^{q_{n+1}}(\Gamma)$ with the estimate $\|(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\|_{0,q_{n+1};\Gamma}\leq\|\nabla_{\Gamma}\mathbf{u}\|_{0,\frac{dq_{n}}{d-q_{n}};\Gamma}\|\mathbf{u}\|_{0,\frac{dq_{n}}{d-2q_{n}};\Gamma}\lesssim\|\mathbf{u}\|_{2,q_{n};\Gamma}^{2}$. Therefore, by 4.8, we deduce that $(\mathbf{u},\uppi)\in\mathbf{W}^{2,r_{n+1}}_{t}(\Gamma)\times\mathrm{W}^{1,r_{n+1}}_{\#}(\Gamma)$, with $r_{n+1}:=\min\{p,q_{n+1}\}$ and